An athlete throws the ball up to a height of 6 what is the i risk velocity of the tennis ball.
1
Expert's answer
2020-10-27T00:37:40-0400

According to the mechanical energy conservation, the kinetic energy at a time of launch is equal to the potential energy in the highest position:


"\dfrac{mv_i^2}{2} = mgh"

where "m" is the mass of the ball, "v_i" is the initial velocity, "h = 6m" is the height and "g = 9.81m\/s^2" is the acceleration due to gravity. Expressing the intial velocity, obtain:


"v_i = \sqrt{2gh} = \sqrt{2\cdot 9.81\cdot 6} \approx 10.85\space m/s"

Answer. 10.85 m/s.